# Unifying KV Cache Compression for Large Language Models with LeanKV

### v1

@@ 7. Evaluation
We first evaluate the effectiveness of LeanKV’s unified KV cache compression algorithm and analyze the sensitivity of different models to KV cache compression. Then, We demonstrate the improvements in throughput achieved by LeanKV.
7.1. Experiment Setup
We evaluate LeanKV on six models across three families: Llama2-7B and 70B (touvron2023llama, ), Llama3-8B and 70B (dubey2024llama, ), Mistral-7B (jiang2023mistral, ), and Mixtral-8x7B (jiang2024mixtral, ). Mixtral-8x7B employs a Mixture-of-Experts (MOE) architecture, while the remaining models are dense. Performance of LeanKV is evaluated on five benchmarks, grouped into two categories. The first includes tasks requiring advanced reasoning and logic: GSM8K (cobbe2021training, ) (mathematics), HumanEval (chen2021evaluating, ) (coding), and MMLU (hendrycks2020measuring, ) (academic level question answering). The second category includes traditional text processing tasks: SQuAD (rajpurkar2018know, ) (reading comprehension) and CNN-DailyMail (nallapati2016abstractive, ) (summarization). Throughput and latency is measured on Nvidia L40 GPUs, each with 48GB of memory (L40, ).
7.2. Evaluating Unified Cache Compression
@@
We first demonstrate the effectiveness of Hetero-KV quantization and dynamic sparsity individually, and then evaluate the end-to-end benefits of the proposed unified KV cache compression. To mitigate the noise and randomness in floating-point computations, each experiment is repeated five times with the dataset randomly shuffled, and the reported results are the average of these runs.
Evaluating Hetero-KV. Figure 7 shows the performance of Hetero-KV quantization, namely K8V4 and K4V2, compared to FP16 baselines, on GSM8K and SQuAD, representing benchmarks from the two categories. To validate our intuition that key vectors have a more significant impact than value vectors (Section 3.1), we additionally evaluate the mirror configurations of Hetero-KV, namely K4V8 and K2V4, where values are stored with higher precision than keys.
K8V4 maintains the accuracy of the FP16 baseline across both benchmarks for all models, whereas its mirror configuration, K4V8, results in noticeable degradation, particularly on the reasoning-intensive GSM8K benchmark. The performance gap between K4V2 and K2V4 is even more pronounced. On GSM8K, K4V2 retains at least 77.8% of FP16 performance, while K2V4 reduces accuracy to nearly zero for all models. On SQuAD, K4V2 achieves 94.7% to 99.4% of FP16 performance, whereas K2V4 achieves only 28.2% to 60.0%. These results confirm that key vectors play a more critical role than value vectors, and demonstrate that Hetero-KV is effective in maintaining the accuracy of the FP16 baseline.
Figure 7. Performance of Hetero-KV.
Evaluating Dynamic Sparsity. We evaluate the effectiveness of per-head dynamic sparsity (Section 3.2) in identifying critical tokens, comparing it to the static sparsity method used in H2O (zhang2024h2o, ), which allocates an equal memory budget to all attention heads. The results are shown in Figure 8, where we use the Mistral-7B model as an example and all tokens are stored in FP16.
For reasoning-intensive workloads such as GSM8K and HumanEval, dynamic sparsity significantly outperforms static sparsity, with the performance gap widening as the fraction of pruned tokens increases. For traditional text processing tasks, however, the benefits of dynamic sparsity are less pronounced: it slightly outperforms static sparsity on SQuAD and performs similarly on CNN-DailyMail.
The difference between SQuAD and CNN-DailyMail can be attributed to their differing sensitivities to token pruning. SQuAD (reading comprehension) requires precise retrieval of specific words from the prompt, making it more sensitive to pruning critical tokens. In contrast, CNN-DailyMail (summarization) inherently involves discarding less relevant information to generate concise summaries, resulting in lower sensitivity to token pruning. To summarize, dynamic sparsity is superior to static sparsity for reasoning-intensive tasks and text processing that involves precise information retrieval.
Figure 8. Performance of dynamic vs. static sparsity.
Evaluating Unified KV Cache Compression. Finally, we evaluate the end-to-end performance of the unified KV cache compression described in Section 4. For each benchmark, we perform grid search over the high-precision threshold $T_{h}$ (with discrete values from {$0.5\%$, $1\%$, $5\%$, $10\%$, $20\%$}) and the low-precision threshold $T_{l}$ (with discrete values from $0\%$, $0.1\%$, $0.5\%$, $1\%$}), selecting the configurations that minimize the KV cache memory budget under different performance constraints. To account for the inherent randomness in floating-point computations, we consider a configuration to be performance-neutral if its average performance falls within three standard deviations of the FP16 baseline.
Table 1 reports the KV cache memory usage of LeanKV, normalized to vLLM (kwon2023efficient, ), across all evaluated models and benchmarks. The results are presented for two configurations: performance-neutral, which ensures no degradation in output quality, and less-than-$5\%$ performance degradation, a common target in prior work (lee2024infinigen, ; fu2024moa, ; zhao2024atom, ; lin2024qserve, ; lin2024awq, ; frantar2023sparsegpt, ) that enables more aggressive memory savings while preserving acceptable quality. Across models and workloads, LeanKV achieves performance neutral with $18.9\%$ to $28.9\%$ of the KV cache memory budget, and less-than-$5\%$ performance degradation with $7.6\%$ to $20.6\%$ of the memory budget.
Additionally, Figure 9 shows the Pareto performance-memory curves of LeanKV on the reasoning-intensive benchmark GSM8K and text processing benchmark SQuAD. The results are compared against the FP16 baseline, conventional 8-bit quantization (K8V8) and 4-bit quantization (K4V4), and the pruning-only technique H2O (zhang2024h2o, ). For GSM8K, K8V8 preserves the FP16 baseline performance, while K4V4 leads to a noticeable performance degradation of $2.8\%$ to $7.3\%$. In comparison, LeanKV achieves FP16-equivalent performance with only $19.8\%$ to $23.1\%$ of the memory budget, even less than K4V4, which requires $25\%$ memory. Moreover, LeanKV offers a more flexible memory-performance tradeoff, with performance degradation limited to less than $5\%$ while using just $9.4\%$ to $16.3\%$ of the memory. In contrast, under similar memory constraints, H2O results in near-zero accuracy. Similar trends are observed on the SQuAD benchmark: LeanKV matches the performance of FP16 and K8V8 with less memory than K4V4, which causes noticeable performance degradation, and outperforms H2O under comparable memory budgets.
Notably, across all evaluated models and benchmarks, LeanKV achieves performance-neutral with the K8V4-K4V2 configuration, where critical tokens are stored with K8V4 precision, non-critical tokens with K4V2, and no pruning is applied. For the less-than-$5\%$ degradation setting, the optimal configuration varies across models. For large 70B models, less-than-$5\%$ degradation is achieved with the more aggressive K4V2-K4V1 quantization without pruning, as pruning significantly degrades performance even on top of K8V4-K4V2. In contrast, for smaller models, less-than-$5\%$ degradation is achieved by further pruning the least important tokens on top of K8V4-K4V2. These varying configurations in low-memory regimes are also visualized in Figure 9, where each data point for LeanKV is annotated with its corresponding configuration. This disparity underscores the varying sensitivity of models to KV cache compression techniques, emphasizing the need for a unified framework capable of adapting to diverse scenarios to maximize efficiency.
In summary, LeanKV provides a superior performance-memory tradeoff compared to conventional methods. It maintains the performance of FP16 and 8-bit quantization while using even less memory than 4-bit quantization. Additionally, LeanKV offers a smooth tradeoff in the low-memory regime ($10\%$ to $20\%$ of memory), where pruning-only methods experience significant performance degradation.
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
@@
Figure 9. KV cache memory normalized to vLLM vs. benchmark performance tradeoff of LeanKV
@@
Table 1. Memory usage of LeanKV normalized to vLLM across models and benchmarks.
Models Benchmarks GSM8K MMLU HumanEval SQUAD CNN-DailyMail Llama2-7B perf-neutral 18.9% 22.8% 23.1% 21.2% 21.6% 5% perf-loss 11.3% 11.5% 18.9% 14.6% 7.6% Llama2-70B perf-neutral 21.3% 21.4% 21.2% 21.1% 20.1% 5% perf-loss 19.5% 15.9% 18.2% 16.0% 15.7% Llama3-8B perf-neutral 20.7% 26.7% 26.6% 24.1% 20.0% 5% perf-loss 9.4% 20.6% 22.1% 12.8% 8.8% Llama3-70B perf-neutral 23.1% 21.5% 22.7% 22.7% 19.5% 5% perf-loss 16.3% 16.1% 21.2% 16.0% 15.7% Mistral-7B perf-neutral 26.8% 28.9% 23.0% 26.2% 19.4% 5% perf-loss 17.7% 13.6% 15.1% 15.9% 9.7% Mixtral-8x7B perf-neutral 19.8% 20.1% 21.4% 21.1% 20.0% 5% perf-loss 13.8% 8.9% 10.7% 8.9% 10.0%
7.3. Sensitivity to KV Cache Compression
We investigate the varying sensitivity of models to KV cache compression, using the Llama model family as an example, due to their similar training data and methodologies. Figure 10 illustrates the normalized performance degradation relative to the FP16 baseline for Llama2-7B, Llama3-8B, and Llama3-70B as the fraction of pruned tokens increases. Among the models, Llama2-7B exhibits the least performance degradation, even when compared to Llama3-8B, which has a comparable parameter size. This resilience can be attributed to Llama2-7B’s MHA architecture, which introduces greater redundancy in the KV cache compared to the more memory efficient GQA architecture utilized by Llama3-8B and Llama3-70B. Llama3-70B experiences the most significant performance degradation, suggesting that larger models distribute attention more evenly across tokens, making them more sensitive to pruning.
Figure 11 further illustrates the normalized performance degradation of the K4V2-K4V1 configuration as the fraction of tokens stored in K4V1 increases. The Llama3-70B model exhibits the greatest resilience to low-bit quantization, maintaining near-baseline performance even when 60% of tokens are stored in K4V1. In comparison, smaller models show noticeable performance degradation even when all tokens are stored in K4V2, highlighting their greater sensitivity to aggressive quantization.
These observations justify the two key trends outlined earlier: larger models and newer models employing more memory-efficient network architectures are more sensitive to pruning, while smaller models are more adversely impacted by low-bit quantization.
Figure 10. Sensitivity to KV cache pruning.
Figure 11. Sensitivity to low-bit KV cache quantization.
7.4. Evaluating Latency and Throughput
@@
In this section, we provide a comprehensive evaluation of LeanKV’s performance, comparing it against state-of-the-art systems including vLLM and H2O across various models and configurations. All experiments were conducted on a machine equipped with eight NVIDIA L40 GPUs. For models that fit into a single GPU’s memory, such as Llama2-7B and Llama3-7B, we use one GPU. For larger models like Llama2-70B and Llama3-70B, we parallelize the computation across four GPUs. Our evaluation focuses on latency breakdown, attention kernel speedup, and end-to-end throughput to demonstrate the efficiency and scalability of LeanKV.
@@
Memory Management Overhead. Figure 12 shows the latency breakdown of LeanKV during one inference step with a sequence length of 1024 tokens, evaluated across different batch sizes. The memory management overhead is remarkably low, contributing less than 0.2% of the total latency in the prompt phase and under 0.9% in the generation phase. Model execution dominates the latency in both phases, accounting for 96–97% of the latency in the prompt phase and 92–93% in the generation phase. Notably, as the batch size increases, the percentage of time spent on model execution rises slightly, reflecting the scalability of LeanKV. For comparison, in vLLM, model execution contributes 97% of the latency in the prompt phase and 95% in the generation phase. These results demonstrate that LeanKV effectively minimizes latency contributions from memory management and KV compression, ensuring that these auxiliary processes do not become performance bottlenecks.
Figure 12. Latency breakdown of LeanKV.
@@
Figure 13. Attention kernel speedup.
End-to-End Throughput. Figure 14 presents the end-to-end throughput and achieved batch sizes of LeanKV compared to vLLM and H2O. We conduct experiments on 1000 sequences sampled from the ShareGPT (ShareGPT, ) dataset, setting the generation length to 4096 tokens. The compression thresholds for LeanKV are selected based on the most challenging dataset, MMLU, making them the most conservative choices. In configurations with no accuracy loss, LeanKV achieves throughput speedups ranging from 2.0×2.0$\times$2.0 × to 2.5×2.5$\times$2.5 × compared to vLLM across various models. When allowing a 5% accuracy loss, LeanKV achieves even higher speedups, ranging from 3.0×3.0$\times$3.0 × to 6.9×6.9$\times$6.9 ×. These improvements are directly correlated with the larger batch sizes supported by LeanKV, enabled by its KV cache compression techniques. For instance, with the Llama2-7B model, LeanKV sustains a batch size of 62 without accuracy loss and 177 with 5% accuracy loss, far exceeding vLLM’s batch size of 22. The increased batch size enables LeanKV to maximize the utilization of GPU compute capacity, leading to higher throughput. Llama3-7B exhibits a less significant speedup compared to Llama2-7B because it utilizes GQA, which reduces its KV cache size by 4$\times$. As a result, vLLM already supports a relatively large batch size of 51.5 for Llama3-7B, narrowing the batch size advantage that LeanKV can leverage. Nevertheless, LeanKV still achieves up to 3.4$\times$ higher throughput over vLLM on Llama3-7B.
Notably, H2O supports even smaller batch sizes than vLLM and consequently has lower throughput. This is because H2O consumes additional memory during the prompt phase for intermediate results required to calculate compression metrics, limiting the batch size that can be supported. In contrast, we design a FlashAttention-compatible kernel using Triton, which avoids the additional memory overhead caused by intermediate results and achieves performance on par with vLLM.
Figure 14. Throughput and achieved batch size.
8. Conclusion
We propose LeanKV, a unified KV cache compression framework designed to enhance LLM serving efficiency through three key techniques: Hetero-KV, per-head dynamic sparsity, and unified KV compression. To support these innovations, LeanKV incorporates system optimizations including unified paging and on-GPU parallel memory management. Our experiments show that LeanKV achieves 1.9×1.9$\times$1.9 × to 2.5×2.5$\times$2.5 × higher throughput over state-of-the-art systems without compromising model accuracy, and up to 6.9×6.9$\times$6.9 × higher throughput with minimal accuracy degradation.

### v2

@@ 7 Evaluation
We first evaluate the effectiveness of LeanKV’s differentiated KV cache compression policy and analyze the sensitivity of different models to KV cache compression. Next, we evaluate the efficiency of the memory manager as well as the end-to-end throughput improvements achieved by LeanKV.
7.1 Experiment Setup
We evaluate LeanKV on five models spanning three major families: Llama3-8B and 70B [60], Qwen2.5-7B and 32B [28], and the recently released thinking model QwQ-32B [58], which generates extended chains of thought and exhibits strong performance on complex reasoning tasks. Our evaluation covers a diverse set of mainstream benchmarks across multiple task domains, including general knowledge (MMLU [26] and MMLU-Pro [55]), mathematics (GSM8K [13] and MATH [34]), code generation (HumanEval+ [11, 39] and MBPP+ [7, 39]), and long-context understanding (LongBench [8]). In addition, we assess LeanKV’s performance on QwQ-32B using two particularly challenging benchmarks: AIME24 [3], a high-school level mathematics competition dataset, and GPQA [48], which evaluates graduate-level science reasoning. Throughput and latency are measured on NVIDIA L40 GPUs, each equipped with 48 GB of memory [45].
7.2 Differentiated KV Compression Policy
@@
We first assess the effectiveness of differentiated KV quantization and dynamic sparsity individually, and then evaluate the combined benefits of the proposed differentiated KV cache compression. To ensure the reliability of our results and mitigate the influence of noise and randomness inherent in floating-point computations, each experiment is repeated five times with the dataset randomly shuffled, and the reported results represent the average across these runs.
Evaluating Differentiated KV Quantization. Figure 7 shows the performance of differentiated KV quantization on GSM8K and HumanEval+, namely K8V4 and K4V2, compared to FP16 baselines, across Llama-3 8B & 70B and Qwen-2.5 7B. To validate our intuition that keys have a more significant impact than values (Section 3.1), we additionally evaluate the mirror configurations, namely K4V8 and K2V4, where values are stored with higher precision than keys. Results show that K8V4 matches the accuracy of the FP16 baseline across all models and benchmarks. In contrast, its mirror configuration, K4V8, exhibits noticeable performance degradation, particularly for Qwen2.5-7B, where accuracy drops to nearly zero on both tasks. This pronounced sensitivity to 4-bit key quantization in Qwen2.5-7B is likely due to its GQA architecture, which applies aggressive KV compression with a queries-per-KV ratio of 7, substantially higher than the ratio of 4 used in Llama-3 8B. Similarly, K4V2 retains over 65% of FP16 accuracy across both benchmarks on Llama-3 8B and 70B, while the mirror configuration K2V4 results in near-zero accuracy. These results confirm that keys play a more critical role than values, and demonstrate the effectiveness of differentiated KV quantization in preserving accuracy while enabling aggressive compression.
Figure 7: Performance of differentiated KV quantization.
Figure 8: Performance of dynamic vs. static sparsity.
Evaluating Dynamic Sparsity. We evaluate the effectiveness of per-head dynamic sparsity (Section 3.3) in identifying critical tokens, comparing it to the static sparsity method used in SnapKV [35] and H2O [69], which allocates an equal memory budget to all attention heads. Figure 8 reports results on Qwen2.5-7B and Llama3-8B across GSM8K and HumanEval+. The x-axis indicates the percentage of tokens pruned, while the y-axis shows the resulting task accuracy. Dynamic sparsity significantly outperforms static sparsity in Llama3-8B, maintaining full accuracy with 50% tokens pruned on GSM8k, and 10% tokens pruned for the more sensitive HumanEval+. For Qwen2.5-7B, dynamic sparsity also consistently surpasses static sparsity, achieving accuracy higher than 80% on GSM8k with 50% tokens pruned. While the improvement is less pronounced on HumanEval+, dynamic sparsity still yields higher accuracy at equivalent pruning ratios. In summary, per-head dynamic sparsity is superior to static sparsity by leveraging head-specific significance.
Figure 9: Calibrating the high- and low-precision threshold parameters $\alpha_{h}$ and $\alpha_{l}$ on the training split of the MATH dataset.
Parameter Calibration. As described in Section 4, LeanKV compares the significance score of a token with its theoretical average, $\frac{1}{N}$, where $N$ denotes the sequence length. A token is quantized to high precision if its score exceeds $\frac{\alpha_{h}}{N}$, to low precision if its score falls within the interval $[\frac{\alpha_{l}}{N},\frac{\alpha_{h}}{N}]$, and is pruned otherwise. To calibrate the appropriate parameters $\alpha_{h}$ and $\alpha_{l}$, we utilize the training split of MATH [34]. We select MATH as the calibration dataset as it is comprised of complex mathematical reasoning tasks with high text information density, where even the removal of a single symbol can invalidate an argument. This makes MATH a highly suitable choice for parameter fitting, as the resulting compression parameters are expected to generalize well across other tasks. Additionally, MATH provides a dedicated training split, ensuring that parameter fitting does not involve testing data and mitigating potential overfitting.
We profile a consistent range of parameters across all evaluated models, as depicted in Figure 9. The x-axis represents the profiled parameter values, while the y-axis shows the resulting accuracy on the calibration dataset. Specifically, we profile $\alpha_{h}$ within the integer range $[1,5]$, progressively relaxing the high-precision threshold beyond the theoretical average. This adjustment accounts for aggregating scores from multiple attention heads by maximum for GQA, as mentioned in Section 4. The profiled values of $\alpha_{h}$ achieve FP16-equivalent accuracy across all models, except for Qwen2.5-7B, which exhibits heightened sensitivity to 4-bit key quantization. Based on our profiling, we select the values of $\alpha_{h}$ that yield the highest accuracy for each model: $\alpha_{h}=1$ for Llama-3 8B and 70B, $\alpha_{h}=3$ for Qwen2.5-32B and QwQ-32B, and $\alpha_{h}=0$ for Qwen2.5-7B. Subsequently, we profile the low-precision threshold $\alpha_{l}$, while keeping $\alpha_{h}$ fixed. We observe that setting $\alpha_{l}=0.1$ induces at least 5% accuracy degradation on the larger models. As a result, we profile five equally spaced values within the range $[0,0.1]$ and select the one yielding the highest accuracy for each model. The final selected values of $\alpha_{l}$ are $0.02$ for Llama3-8B, $0.04$ for Qwen2.5-7B, and 00 for the remaining models. While the profiled values cover a limited portion of the entire parameter space, as will be demonstrated in the following sections, LeanKV significantly outperforms existing baselines with these fitted parameters, highlighting its potential effectiveness.
@@
Figure 10: KV cache memory normalized to vLLM vs. benchmark performance tradeoff of LeanKV.
Evaluating Differentiated Compression Policy. Finally, we evaluate the combined benefits of LeanKV’s KV cache compression policy. We compare LeanKV against several state-of-the-art baselines, including pruning-based methods such as H2O [69], SnapKV [35], and DuoAttention [63], quantization-based approaches such as 4-bit KV and KIVI [41], as well as KV partial loading techniques like Quest [56]. Table 1 summarizes the accuracy and KV cache memory usage of LeanKV normalized to vLLM [32] baseline, with the parameters tuned as described, for the evaluated non-thinking models. For the pruning-based baselines and Quest, we tune their memory usage to 50% of the baseline. LeanKV achieves near-lossless accuracy relative to the FP16 baseline, with an average degradation of only $0.3\%$, while using $19.3\%$ to $36.7\%$ memory. In contrast, all other baselines incur more significant accuracy degradation at comparable or higher memory usage. Furthermore, LeanKV dynamically adapts its memory usage based on the task’s information density. For example, LeanKV allocates less memory for the 5-shot MMLU benchmark compared to the 0-shot HumanEval+, where the prompt consists solely of the function name and comments.
LeanKV also outperforms baseline methods in long-context scenarios, as shown in Table 2, where we present results from one benchmark in each category of LongBench. For Llama3-8B and Qwen2.5-7B, LeanKV achieves memory usage of $14.9\%$ and $27.0\%$, respectively, while baseline methods are tuned to $25\%$ memory usage. Notably, LeanKV consistently achieves superior accuracy across all categories, with an average degradation of only $0.4\%$, while utilizing less or comparable memory compared to the baseline methods.
Furthermore, the performance gap between LeanKV and baseline methods becomes even more pronounced for the thinking model QwQ-32B on more challenging benchmarks, as demonstrated in Table 3. LeanKV matches the FP16 baseline performance using only $27.4\%$ of the memory on average, while all other methods experience significant performance degradation. On the math competition benchmark AIME24, the most accurate baseline, 4-bit KV quantization incurs a $16.0\%$ accuracy degradation. On the graduate-level science benchmark GPQA, the accuracy of pruning-based and KV partial loading methods drops to almost zero.
Lastly, Figure 10 depicts the performance-memory trade-offs of LeanKV across representative benchmarks under the profiled parameter values, namely $\alpha_{h}\in[0,5]$ (step size of 1) for all models and $\alpha_{l}\in[0.02,0.1]$ (step size of 0.02) for Llama3-8B and Qwen2.5-7B. We also include the accuracy of baseline methods within similar memory usage ranges for comparison. LeanKV consistently matches the FP16 baseline accuracy across different models and benchmarks within the profiled range of memory usage, demonstrating its robustness. In contrast, baseline methods experience significant accuracy degradation, particularly pruning-based and KV partial loading methods, which suffer rapid accuracy declines as memory usage decreases. In summary, LeanKV provides a superior performance-memory tradeoff compared to state-of-the-art methods.
@@
Table 1: Accuracy and memory usage of LeanKV and baseline methods across models and benchmarks. Numbers in parentheses represent memory usage normalized to FP16. The memory usage is 50% for SnapKV, Quest, H2O and DuoAttn, 25% for INT4, and 12.5% for KIVI due to 2-bit quantization.
[…]
Table 2: Evaluating LeanKV on LongBench. LeanKV uses 14.9% memory for Llama3.1-8B and 27.0% for Qwen2.5-7B. For Quest and SnapKV, the memory usage is 25%.
[…]
Table 3: Evaluating LeanKV on QwQ-32B thinking model.
[…]
7.3 System Performance
@@
In this section, we provide a comprehensive evaluation of LeanKV’s performance, comparing it against state-of-the-art systems including vLLM, Quest, and SnapKV across various models. All experiments are conducted on a machine equipped with eight NVIDIA L40 GPUs. For models that fit within the memory of a single GPU, such as Llama3-8B and Qwen2.5-7B, we use one GPU. For larger models, we parallelize the computation across multiple GPUs: four GPUs for Llama3-70B, and two GPUs for Qwen2.5-32B and QwQ-32B. Our evaluation focuses on latency breakdown, attention kernel speedup, and end-to-end throughput to demonstrate the efficiency and scalability of LeanKV.
Memory Management Overhead. Figure 11 compares LeanKV’s on-GPU parallel KV compaction with an alternative implementation that performs multi-threaded memory management on the CPU. Evaluated with a sequence length of 1024 tokens across different batch sizes, LeanKV reduces memory management latency by up to three orders of magnitude relative to the on-CPU approach. For one entire inference step, LeanKV significantly outperforms the on-CPU approach, particularly in the generation phase, where the memory management overhead on the CPU far exceeds the model execution time. Therefore, LeanKV’s on-GPU parallel memory management is essential for ensuring that the performance benefits of KV cache compression are fully realized without being overshadowed by memory management overhead.
@@
Figure 12 further illustrates the latency breakdown of LeanKV during one inference step. The memory management overhead is remarkably low thanks to LeanKV’s on-GPU parallel KV compaction, contributing less than 0.2% of the total latency in the prompt phase and under 0.9% in the generation phase. Model execution dominates the latency in both phases, accounting for 96–97% of the latency in the prompt phase and 92–93% in the generation phase. Notably, as the batch size increases, the percentage of time spent on model execution rises slightly, reflecting the scalability of LeanKV.
(a) Memory management
(b) One entire inference step
Figure 11: Latency comparison between parallel KV compaction and on-CPU multi-threaded memory management.
Figure 12: Latency breakdown of LeanKV.
@@
Figure 13: Attention kernel speedup.
End-to-End Throughput. Figure 14 presents the end-to-end throughput and achieved batch sizes of LeanKV compared to vLLM, Quest, and SnapKV. The maximum generation length is set to 16K tokens for QwQ-32B, 8K for Qwen2.5-32B, and 4K for the other models. We evaluate on 1000 sequences sampled from the MATH dataset [34], which naturally elicits chain-of-thought reasoning and typically leads to long generations reaching the specified limit. Compression thresholds for LeanKV are adopted from Figure 9. Across all models, LeanKV consistently achieves higher throughput than prior systems. Notably, LeanKV achieves a remarkable 5.4×5.4$\times$5.4 × higher throughput over vLLM on the QwQ-32B thinking model, while Quest and SnapKV achieve only 1.6×1.6$\times$1.6 × and 1.8×1.8$\times$1.8 × speedups, respectively. This result highlights the superior efficiency of LeanKV in processing long sequences, making it particularly well-suited for tasks that require extended reasoning.
The throughput improvements are directly correlated with the larger batch sizes supported by LeanKV, enabled by its KV cache compression techniques. For instance, with the QwQ-32B model, LeanKV sustains a batch size of 15.9, far exceeding vLLM’s batch size of 2.7. The increased batch size enables LeanKV to maximize the utilization of GPU compute capacity, leading to higher throughput. In contrast, Quest supports the same batch size as vLLM because it retains the entire KV cache without reducing memory usage. Its speedup comes primarily from faster attention computation, as it processes only a subset of tokens deemed important. However, the process of estimating token importance incurs additional overhead, limiting the overall efficiency of Quest.
Figure 14: Throughput and achieved batch size.
8 Conclusion
We propose LeanKV, a framework to enhance LLM serving efficiency by exploiting three levels of differentiation in KV cache: differentiated precision for keys and values, hierarchical compression based on varying token importance, and per-head dynamic sparsity. At the core of LeanKV is the parallel KV compaction technique that efficiently handles irregular memory requirements across requests and attention heads, effectively translating memory savings into performance gains. Our evaluation shows that LeanKV is able to compress the KV cache by 2.7×2.7$\times$2.7 × to 5.7×5.7$\times$5.7 × without compromising accuracy on complex workloads that require sophisticated reasoning and long-generation capabilities, and enhances throughput by 1.9×1.9$\times$1.9 × to 5.4×5.4$\times$5.4 ×.